OpenStudy (anonymous):

solve f(x)=0 for x, f(x)=3x^2- 12.

OpenStudy (anonymous):

f(x)=3x^2- 12 0 = 3x^2 - 12 +12 +12 12 = 3x^2 /3 /3 4 = x^2 \[\sqrt{4}= \sqrt{x^2}\]\[2=x\] Does that make sense? :o)

OpenStudy (anonymous):

Well, it'd be 2 and -2 http://www.wolframalpha.com/input/?i=f%280%29%3D3x%5E2-+12 Can you see the graph?